OpenStudy (anonymous):

William is 12 years older than Ben. Nineteen years ago, William was 3 times as old as Ben.How old is Ben now?

OpenStudy (anonymous):

Set up equations.\[W = B+12\]\[W-19 = 3B - 19\]Now solve for B.

OpenStudy (anonymous):

so my equation would be B+12-19=3B-19?

OpenStudy (anonymous):

Correct. I'm not sure if I am right, though. According to equations, Ben doesn't exist 19 years ago...

OpenStudy (anonymous):

Bump your question to get more helps.

OpenStudy (anonymous):

oh okay thanks :]

OpenStudy (fellowroot):

you need to bump harder